February 5th 2021

From Hall A Wiki
Jump to: navigation, search
  • Updates
    • HCal
      • started setting up summing trigger for cosmics, need more discriminators. Discriminator setup, need remote control of threshold ( check with electronics group ) - need threshold control - DAC VME board to be setup ( Chris board ) - need a long cable for setup in the hall. Need evaluate threshold for individual PMTs ( need to think of a scheme ) - could try turning some HV off - capacitive coupling ? ( need to figure out - most likely no) - need have trigger energy calibration - summing module sum of 16 - is 60 Hz an issue ? Added capacitor coupling on BigBite (20 pF)might need to do that ( but AC coupling has rate dependence ) - channel going in discriminator - 80 group - plan to have 1 more FADC for reference time - if 2 FADCs can add sum.
      • HCAL DAQ upgrade to Centos 7 and last FADC library version will take a few days
    • Bigbite shower
      • VME crate setup : took data yesterday
    • Bigbite hodoscope
      • taking runs every day - ADC and TDC spectra - was stopping at 1 M - run in VNC ok - debugging issues - s
    • LHRS
      • all fastbus crate working stand alone
      • will put all crates together
      • still need to install Centos 7
    • INFN GEM
      • taking cosmics since Wednesday - some crashes from time to time - DAQ is fine and low level plots look good - but cosmics data missing large amount
      • Bryan fixed hv program but most likely not related to previous issue
    • UVA GEM clean room
      • still have network issues - router ordered should arrive today
    • GEM TEDF
      • got working common noise subtraction and pedestal subtraction
      • cosmics trigger setup, Sean looking at adjusting latency
    • CDet
      • getting back running - missing hardware for PMTs holders

, Peter will work on installing PMTs

      • clrlpc is back : will be used for CDET DAQ
      • need find TI SD and CPU and waiting for VETROCs
    • GRINCH
      • sealed with gasket, trying powering to check if everything is work - ribbon cable working -

noisy on some cables - around translator

      • CODA3 working, taking data - need to adapt decoder to look at data
    • GEn RP
      • putting detector together

  • For next week : FADC list and priority FADC wish list